OpenStudy (anonymous):

Kenny goes shopping for school clothes during a tax−free weekend, and he saves 6% sales tax. Use a percent proportion to determine how much he would save if he bought clothes $300 worth . Which of the following is true?

OpenStudy (anonymous):

The solution is $18, and is a denominator in the proportion. The solution is $18, and is a numerator in the proportion. The solution is $5,000, and is a denominator in the proportion. The solution is $5,000, and is a numerator in the proportion.

OpenStudy (anonymous):

?

hero (hero):

x/300 = 6/100

hero (hero):

The answer is x = 18 so the first answer choice

OpenStudy (anonymous):

ohh thanks lol

hero (hero):

By the way, the general formula for these is is/of = %/100

hero (hero):

\[\frac{is}{of} = \frac{\%}{100} \]

hero (hero):

Oh wait, lol, it's the second choice

OpenStudy (anonymous):

>:[]

hero (hero):

? Is that a happy face or a sad face?